how do i factor by using the trial and error method for this problem? Help
15x^2+13x+2

Answer by venugopalramana(3286)

15x^2+13x+2
MULTIPLY COEFFICIENTS OF X^2(15) AND CONSTANT TERM (2)..WE GET 15*2=30...FIND 2 FACTORS FOR THIS BY TRIAL AND ERROR SO THAT THEIR SUM IS COEFFICIENT OF X (13)..
NOW FACTORS OF 30 ARE
1,30...SUM=31
2,15...SUM=17
3,10...SUM=13....DONE
SO SPLIT 13X AS 10X AND 3X
15X^2+10X+3X+2=5X(3X+2)+1(3X+2)=(3X+2)(5X+1)